OpenStudy (anonymous):

For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 2x + y = 8 -x - 3y = -12 3y + 12 -3y + 12 3y - 12 -3y - 12

OpenStudy (anonymous):

Why don't you try isolating x in the second equation?

OpenStudy (anonymous):

how would i do that?

OpenStudy (anonymous):

Remember your algebra. You need to undo things that get done to x by doing the opposite to both sides. I'll give you an example, okay? 2x+6y = 10 6y gets added to x, so I need to undo that by doing the opposite. What is the opposite of adding 6y? Subtracting 6y. So I do that to both sides and get: 2x +6y -6y = 10-6y 2x = 10-6y Now, I look and see that x is multiplied by 2. To undo that, I can divide by 2, but I need to do that to both sides so I get: (2x)/2 = (10/2) - (6y/2) x = 5-3y